Alert.png The wiki is deprecated and due to be decommissioned by the end of September 2022.
The content is being migrated to other supports, new updates will be ignored and lost.
If needed you can get in touch with EGI SDIS team using operations @

NGI DE CH Operations Center:Operations Meeting:02122011

From EGIWiki
Jump to navigation Jump to search

Operations Meeting Main


  • Minutes of last meeting


  • Meetings/conferences
NGI-DE/NGI-CH/D-Grid Workshop in April
Note: There is also a dCache Workshop in April. Date should be chosen carefully.
The EGI Community Forum ( will be in Munich 26-30th March 2012 and held in conjunction with the 2nd EMI 
Technical Conference. Abstract submission was open until 2/12/11.
  • Availability/reliability statistics
recomputation done
  • Monitoring
  • Staged rollout/updates

other topics

EMI release / possible infosys errors (UNI-SIEGEN)

Sites with CRYTICAL gstat status: wuppertalprod Uni-Bonn DESY-HH SCAI MaiGRID LRZ-LMU

Round the sites

  • BMRZ-FRANKFURT (Uni Frankfurt)
we updated all our wn's to torque 2.5.7-2 (glite-WN-version-3.2.12-1) and this works fine with the old torque server (2.3.13-1).  
Server we didn't update because of the problem with memory in new version.
This week update of dcache-cms instance to 1.9.12-13 was done
  • FZJuelich
  • Goegrid
  • GSI
  • ITWM (Martin Braun)
  • KIT (GridKa, FZK-LCG2, Dimitri Nilsen, Tobias Koenig)
gLexec updated at WNs
roled based mapping for glexec was requested by atlas
WMS disk full: Problems with ngi-de-nagios portal
  • KIT (Uni Karlsruhe)
  • LRZ
  • MPI-K
  • MPPMU (Cesare Delle Fratte)
   - DOWNTIME 29/11 01/12 dcache upgrade from 1.9.5-28 to 1.9.12-13
   - problems with gridftp doors (solved by Java jdk update to latest packages)
   - dCache: number of movers was increased
   - updated one of the two CREAM
   - installed security fix on Apel box
   - strange lfc failures caused by gpfs partition problems
  • RWTH Aachen
  • SCAI
  • Uni Bonn
services online
  • Uni Dortmund
  • Uni Dresden (Ralph Mueller-Pfefferkorn)
- since about two months problem with our file system, especially with the central nfs file system. The nfs system becomes 
  overloaded. 100s of jobs with 100s of files.
  Paolo/CSCS: We had the same problems. It was fixed by changing the CREAM grubber and we went from Lustre to gpfs and SSD disks 
  for the metadata and for the inode's table.
  • Uni Freiburg (Anton Gamel)
- problems with gsi ssh -> increased movers
- installed additional dCache servers
  • Uni Mainz-Maigrid
  • Uni Siegen
  • Uni Wuppertal
  • CSCS (Paolo)
- maintenance two days ago: firmware update of the disks, lost 4 disks/CMS pool (in contact with CMS)
- test CERNVMFS in preproduction
  • PSI
  • Switch

Note: please update your entry at if needed.

Status ROD

LRZ from 02.2012. 2*2 Shifts
ROD Newsletter Nov. 2011
tickets were not mentioned within 10 days. Be aware of the ROD statistics.
Please pay attention to the Escalation Procedures


If you have additional topics to be discussed during the meeting, please submit them in advance via our email list email list.